The number of people arriving for treatment at an emergency room can be modeled by a Poisson process with a rate parameter of four per hour.
(a) What is the probability that exactly three arrivals occur during a particular hour? (Round your answer to three decimal places.)


(b) What is the probability that at least three people arrive during a particular hour? (Round your answer to three decimal places.)


(c) How many people do you expect to arrive during a 30-min period?
people
Answer #1
Concepts and reason

The concepts used to solve this problem is application of Poisson distribution.

The occurrence of the event within a fixed interval of time with a known mean rate and independently of the occurrence of the last event is called the Poisson distribution.

Poisson distribution has following conditions,

First is nn \to \infty that is the number of trials is indefinite.

Second is p0p \to 0 represents the probability of success is small for each trial.

Third is np=λnp = \lambda which is finite.

Where, λ\lambda is a positive real number.

Fundamentals

For a random variable X, the probability density function of Poisson distribution can be denoted as:

f(x)=eλλxx!f\left( x \right) = \frac{{{e^{ - \lambda }}{\lambda ^x}}}{{x!}}

Where, λisthemeanorexpectedvalueofX\lambda {\rm{ is the mean or expected value of }}X

(a)

Consider X to be a random variable denoting the arrival of people in the room.

As per the provided information the number of people arriving for treatment at an emergency room follows Poisson distribution with a rate parameter of four per hour. So, the random variable X will follow Poisson distribution with the parameter λ\lambda as 4.

That is,

XP(4)X \sim P\left( 4 \right)

The probability of exactly three people arriving during a particular hour can be calculated by substituting the value of X as 3 in the probability density function of Poisson distribution.

P(X=3)=eλλxx!=e4(4)33!=0.195\begin{array}{c}\\P\left( {X = 3} \right) = \frac{{{e^{ - \lambda }}{\lambda ^x}}}{{x!}}\\\\ = \frac{{{e^{ - 4}}{{\left( 4 \right)}^3}}}{{3!}}\\\\ = 0.195\\\end{array}

(b)

The probability of at least 3 people arriving in the room is calculated as,

P(x3)=1P(x2)=1[P(x=0)+P(x=1)+P(x=2)]=1(eλλxx!+eλλxx!+eλλxx!)=1(e4(4)00!+e4(4)11!+e4(4)22!)\begin{array}{c}\\P\left( {x \ge 3} \right) = 1 - P\left( {x \le 2} \right)\\\\ = 1 - \left[ {P\left( {x = 0} \right) + P\left( {x = 1} \right) + P\left( {x = 2} \right)} \right]\\\\ = 1 - \left( {\frac{{{e^{ - \lambda }}{\lambda ^x}}}{{x!}} + \frac{{{e^{ - \lambda }}{\lambda ^x}}}{{x!}} + \frac{{{e^{ - \lambda }}{\lambda ^x}}}{{x!}}} \right)\\\\ = 1 - \left( {\frac{{{e^{ - 4}}{{\left( 4 \right)}^0}}}{{0!}} + \frac{{{e^{ - 4}}{{\left( 4 \right)}^1}}}{{1!}} + \frac{{{e^{ - 4}}{{\left( 4 \right)}^2}}}{{2!}}} \right)\\\end{array}

=1(0.018+0.073+0.147)=0.762\begin{array}{l}\\ = 1 - \left( {0.018 + 0.073 + 0.147} \right)\\\\ = 0.762\\\end{array}

(c)

By the property of Poisson distribution, it can be denoted that mean or expected number of a random variable X following Poisson distribution is equal to the parameter λ\lambda .

Therefore,

MeanorE(X)=λ{\rm{Mean or }}E\left( X \right) = \lambda

As per the provided information the value of parameter λ\lambda is 4, that is, the excepted number of people arriving in the room in an hour is 4. Therefore, the expected number of people arriving in 30 minutes is the half of the mean or λ\lambda . So,

(Expectednumberofpeoplearrivingintheroomin30mins)=(Expectednumberofpeoplearrivingintheroomin1hr)2=λ2=42=2\begin{array}{c}\\\left( \begin{array}{l}\\{\rm{Expected number of }}\\\\{\rm{people arriving in the}}\\\\{\rm{ room in 30 mins}}\\\end{array} \right) = \frac{{\left( \begin{array}{c}\\{\rm{Expected number of people }}\\\\{\rm{arriving in the room in 1 hr }}\\\end{array} \right)}}{2}\\\\ = \frac{\lambda }{2}\\\\ = \frac{4}{2}\\\\ = 2\\\end{array}

Ans: Part a

The probability of exactly three people arriving during a particular hour is 0.195.

Part b

The probability of at least three people arriving during a particular hour is 0.762.

Part c

The expected number of people arriving in 30 minutes is 2.